If you invest $2,000 at the end of each year for five years and you earn 7% interest compounded annually, how much will you have accumulated at the end of the fifth year?
Answer #1

SOLUTION :


Annual interest rate , r = 7% = 0.07

=> (1 + r) = 1.07 


Period, n = 5 years


Annuity amount, A = 2000 ($)


It is ordinary annuity as it is paid at the end of each year for 5 years.


Now, we use the following formula for calculating FV.


FV = A ((1+r)^n - 1)/ r


So, in the given case,


Amount accumulated at the end of 5 years

= FV

= 2000(1.07^5 - 1)/0.07 

= 11501.48 ($) (ANSWER)

answered by: Tulsiram Garg